### v3

@@ Abstract
Analytic Translation Quality Evaluation (TQE), based on Multidimensional Quality Metrics (MQM), has traditionally used a linear error-to-penalty scale calibrated to a reference sample size of 1000100020002000 words. However, linear extrapolation is known to bias judgment on samples of a different size than the reference, over-penalizing short samples and under-penalizing long ones, resulting in persistent misalignment with expert intuition. Building on our previously introduced Multi-Range framework, this paper presents a calibrated, non-linear scoring model that aims to better reflect how human content consumers perceive translation quality across samples of variable length. We present empirical data from three large-scale enterprise environments and shows that acceptable error counts grow logarithmically—not linearly—with sample size. This premise is supported by psychophysical and cognitive evidence, including the Weber–Fechner law and Cognitive Load Theory, which explain why the impact of additional errors diminishes perceptually but grows cognitively with scale. We propose a two‑parameter model that is anchored to a reference tolerance and calibrated from two tolerance points via a one-dimensional root-finding step. The model yields an explicit interval (in closed form) within which the linear approximation stays within $\pm 20\%$ relative error, and it drops into existing evaluation workflows unchanged apart from a dynamic tolerance function. It improves interpretability, fairness, and Inter-Rater Reliability (IRR) across both human and AI-generated translations. By reinforcing and operationalizing a perceptually valid scoring paradigm, this work advances the field toward a more accurate, scalable, and cognitively grounded translation quality evaluation. The proposed model also aims to provide a better scoring model for further AI-based document-level evaluation, congruent with human judgment.
@@
Translation Quality Evaluation, Statistical Quality Measurement, Translation Evaluation Model, Non-Linear Scoring, Empirical-Based Design
@@ 1 Introduction
Analytic Translation Quality Evaluation (TQE) is widely recognized as the gold standard method for assessing translation quality with high precision. It relies on meticulous annotation of linguistic and technical errors using a consistent typology, often derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014, GladkoffEtAl2024], with proper training of annotators. Once the annotations are in place, a scoring model is applied to produce a quantifiable quality assessment. In practical terms, most industry scoring models operate under a linear assumption: the number of allowed errors scales proportionally with the word count of the evaluated sample. This model typically normalizes errors per 1000 or 2000 words to compare them against a user-defined threshold, e.g., no more than five minor errors per 1000 words. The normalized error density is then used to make a pass-or-fail decision or compute a quality score  [^1] Logrus Global Statistics https://logrusglobal.substack.com/archive . This method works reasonably well for evaluation sample sizes close to the reference size—that is, the standard word count used as a baseline for comparing results (typically 1,000–2,000 words). However, problems arise when evaluating smaller or larger samples [gladkoff2021].
Industrial users, practitioners, and researchers have repeatedly reported a mismatch between the linear model and their holistic judgment of translation quality. Waddington’s comparative studies [waddington1999, waddington2001] show that additive, points-off models alone do not align with holistic judgments; combining analytic and holistic perspectives improves reliability. This suggests that perceived translation quality is not a linear function. Evidence from other domains points in the same direction. In speech and video quality, international standards report top-end ‘compression’ in opinion scores and therefore recommend comparison tasks and non-linear mappings (typically logistic) from objective indices to MOS [ituP800, streijl2016, vqeglogistic]. Within translation industry, QA frameworks encode non-linearity explicitly via severity multipliers and critical ‘showstopper’ errors—i.e., one serious defect can outweigh many minor ones [mqm2014, saeJ2450]. Large-scale MT evaluations similarly show that expert, context-aware severity judgments yield system rankings that diverge from simple averages, supporting non-additive scoring [freitag2021]. In usability engineering, a single ‘catastrophic’ issue overrides numerous small problems, mirroring stakeholders’ acceptance behavior for translations [nielsen1994]. Together, these applied literature reinforce the conclusion that linear points-off schemes understate how overall quality is perceived and acted upon.
Despite its known practical shortcomings, the assumption of linear error scaling remains entrenched in industry practice—partly due to implicit conventions and the replication of established procedures, and partly because people often subconsciously apply linear extrapolation to numerical judgments, including those about error tolerance. Yet when evaluators (and therefore also content users) assess samples without relying on mental calculation, their judgments reliably deviate from linear expectations. [^2] We list some real-world institutional users’ experiences on translation evaluation in Appendix A.
This paper presents empirical evidence and theoretical justification for a non-linear scoring model, first proposed in our earlier work on Multi-Range TQE [GladkoffEtAl2024]—one that more accurately captures how content users perceive translation quality as a function of sample length. The present study argues that as the evaluated text gets longer, the maximum number of allowable errors grows, but more and more slowly—equivalently, the perceptionally allowed errors per 1,000 words decline as word count increases. We capture this pattern with a logarithmic dependency. The Multi-Range TQE work introduced a logarithmic scorecard. This paper shows where that curve comes from: tolerance grows sublinearly with length, as demonstrated by expert surveys and by well-known psychophysical and cognitive results (Weber–Fechner; Cognitive Load Theory). We also provide a simple, easily calibrated model for use in current evaluation tools. The structure of this paper is organized as below: Section 2 introduces the background on analytical TQE and MQM, Section 3 presents our empirical evidence of non-linearity, Section 4 offers psychophysical and cognitive foundation, Section 5 details the non-linear scoring model, Section 6 and 7 cover the practical implications and step-by-step guide on building a non-linear scorecard, followed by Limitations and Future Work (Section 8).
@@ 2.1 Analytic TQE and MQM
Analytic TQE refers to a structured method of assessing translation output by identifying and categorizing errors in reference to the source text and applicable specifications. Unlike holistic methods, which rely on a general impression of quality, analytic TQE requires systematic annotation of specific error types, each classified under predefined categories. The most prominent and widely adopted framework for such structured annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014]. MQM provides a hierarchical error typology that is flexible and extensible, with high-level error types like Accuracy and Linguistic Conventions (formerly called Fluency) and more granular error subtypes such as Mistranslation, Omission, Grammar, and Spelling. Many specialized metrics used across the industry are subsets or derivatives of MQM, often customized to the type of content or the user’s preference.
Once the annotation phase is complete, a scoring model is applied to an evaluation sample (a defined word count from the job), converting the annotated errors into a numeric score or a Pass/Fail verdict. Every MQM-based metric has two inseparable parts: (i) an error typology that specifies which error types are annotated (with a severity level such as minor/major/critical), and (ii) a scoring model that maps those annotations to severity-weighted penalty points and a quality score. For comparability across jobs of different lengths, results are expressed against a reference size—typically 1,000–2,000 words  [^3] https://themqm.org/.
@@ 2.2 Linear Scoring Models and Their Limits
The dominant approach to scoring in translation quality evaluation is the linear model. Here, the severity-weighted penalty total (APT, Absolute Penalty Total—the sum of points assigned to annotated errors) is scaled to a reference sample size so that assessments are comparable across jobs of different lengths. For example, a limit of five minor errors one point each per 1,000 words implies 2.5 penalty points in a 500-word sample and 10 in a 2,000-word sample (and if a major error carries a $5\times$ severity multiplier, a single major error consumes the full 5-point allowance for a sample of 1000 words).
Although this proportional scaling appears objective, it fails to reflect how human content users actually perceive quality. Readers are usually stricter with short texts, where even a single error stands out, and more lenient with longer texts [waddington1999, waddington2001, mqm2014, saeJ2450]. However, this tolerance wanes as the document length grows further and the errors begin to compound. For instance, as noted above, a single major error is often sufficient to fail a one-page sample; in a seven-page sample, however, evaluators would fail it at far fewer than seven major errors—even though a linear rule would imply proportional tolerance. Such intuitive resistance reveals that human perception of quality is fundamentally non-linear.
@@
Building on our earlier work on Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three sample size regimes:
  • micro-range (<250<250 words): For very short samples (under 250 words), traditional quality scoring methods are unreliable. As shown in [gladkoff2021], quality scores for such brief samples exhibit extremely wide confidence intervals—meaning the range of possible true scores is so large that the measurement is imprecise—making individual measurements meaningless for decision-making.
    This limitation is particularly important for automatic Quality Estimation (QE), especially with the practice of sentence- or segment-level scoring. Recent industry findings confirm that these fine-grained predictions do not align with how humans perceive quality. Bane [baneATA2025] highlighted the repeated failures of sentence- or segment-level QE in high-volume translation, where automated scores fail to identify all major, and even critical, errors, making the method unreliable. As discussed in [gladkoff2025substack], attempting to derive meaningful scores from very small samples or isolated sentences or segments simply does not work.
    Instead, for micro-range samples, Statistical Quality Control (SQC) methods offer a more statistically robust and risk-aware framework for interpreting quality signals, explicitly framing the producer’s and consumer’s risk associated with very limited observations.
@@
  • meso-range ($\approx 250\!-\!3{,}000$ words): While a Raw score can be useful for certain scenarios, the recommended approach for this range is a Linear Calibrated Scoring Model, anchored to a user-defined tolerance in severity-weighted penalty points (APT) for a reference wordcount.
@@
  • macro-range (>3,000>3{,}000 words): Human perception diverges from linear scaling calibrated to a reference sample size in the meso-range, so another model is needed: a non-linear (logarithmic) tolerance curve of a Non-Linear Calibrated Scoring Model can instead cover both the meso-range and the macro-range without requiring multiple linear approximations.
@@
The present study concentrates on the entire span of statistically reliable measurements in both the meso-range and the macro-range and supplies both psychophysical and empirical evidence for adopting a logarithmic scoring model across these ranges. Here “statistically reliable” is meant in the technical sense—i.e., estimates with reasonably narrow confidence intervals —the usual sampling-uncertainty bands around an estimate that shrink as sample size grows. [^4] For intuition, a simple 95% Wald interval for a proportion is $p\pm 1.96\sqrt{p(1-p)/n}$ [gladkoff2021]. For small samples or extreme proportions, this estimate is known to undercover; a Wilson (score) interval—or its closely related Agresti–Coull “plus-four” approximation—provides markedly better coverage [AgrestiCoull1998]. Wilson 95% CI: $\frac{\hat{p}+\frac{z^{2}}{2n}\pm z\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}}{4n^{2}}}}{1+\frac{z^{2}}{n}},$ with z=1.96z\!=\!1.96; the Agresti–Coull “plus-four” uses $\tilde{p}=(x+\tfrac{z^{2}}{2})/(n+z^{2})$ and $\tilde{p}\pm z\sqrt{\tfrac{\tilde{p}(1-\tilde{p})}{n+z^{2}}}$ [AgrestiCoull1998].
@@
When linear scoring is applied indiscriminately across all ranges, it produces systematic friction. Users and linguists must override or reinterpret scores, or use several different linear models, especially for very short or very long samples. Some organizations respond by limiting evaluations to a fixed 1,0001{,}000-word sample, but this workaround is operationally restrictive and fails to address the underlying perceptual misalignment. Because the width of a confidence interval shrinks roughly like $1/\sqrt{n}$—with an additional finite-population correction $\sqrt{(N-n)/(N-1)}$ when the evaluated sample is a large share of the document—variation in the sampled word count can change those intervals considerably. [^5] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
The persistent use of linear models—despite long-recognized shortcomings—stems largely from the absence of simple, theoretically grounded alternatives. This paper aims to fill that gap by presenting a non-linear scoring model and providing empirical and theoretical support for it, grounded in psychophysics and cognitive-load research.
@@ 3 Empirical Evidence of Non-Linearity
To investigate how human perception of translation quality varies with sample length, an empirical study was conducted with quality managers from three large institutional users (anonymized). Each of these organizations manages multilingual translation programs with substantial daily volumes and relies on mature, metric-based TQE systems.
@@ 3.1 Survey Design
The extended calibration questionnaire described in [GladkoffEtAl2024] was reused, asking survey respondents to state acceptable error counts for multiple evaluation sample sizes without performing linear extrapolation. This method again proved effective in eliciting a latent non-linear response curve.
A detailed calibration questionnaire was designed to capture Quality Managers’ intuition about acceptable error tolerance—operationalized as the maximum number of severity-weighted penalty points (APT) they would still pass—across a range of evaluation sample sizes.
Importantly, specific measures were taken to neutralize the influence of the standard linear model, which respondents are accustomed to using in their daily work. Without intervention, professionals often default to mental linear math—extrapolating tolerances from known thresholds. To counteract this bias, the institutions’ Quality Managers (i.e., the buying organizations’ quality leads) were asked not to calculate or estimate based on existing formulas, but instead to reflect on what their organization would truly consider acceptable or unacceptable based purely on professional judgment.
The questionnaire asked the following questions:
Intuitive Error Tolerance Questions Intuitively, how many minor errors would you accept in a 1-page sample? Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages? For a 3-page text, how many minor errors seem tolerable—just based on your experience? Do not calculate—just estimate: how many minor errors would you allow in 4 pages? Based on gut feeling, how many minor errors would still be okay in a 5-page sample? Without applying your formal model, what number of minor errors seems acceptable in 10 pages? Without referencing formulas, what minor error count would intuitively feel acceptable for 20 pages?
@@
To reduce linear anchoring, page sizes were presented in randomized order without showing prior answers; two page sizes were repeated at the end to probe intra‑rater stability. Respondents were instructed explicitly not to prorate from known thresholds and to answer from professional judgment alone.
The questionnaire deliberately focused on minor errors for two key reasons. First, higher-severity errors (major and critical) are too infrequent for reliable statistical analysis. Second, as many studies note, their impact on quality perception is significantly different and more pronounced. This specific focus is practical, as standard industry practice allows the severity of any error to be expressed as an equivalent number of minor errors.
While institutions used either words or pages, we normalized all responses for comparison by converting them to a common scale based on a standard of 250 words per page.

### v4

@@ Abstract
Analytic Translation Quality Evaluation (TQE) based on Multidimensional Quality Metrics (MQM) typically applies a linear error-to-penalty mapping calibrated on reference samples of 1000100020002000 words. However, extrapolating this scale to other sample lengths introduces systematic bias, over-penalizing short texts and under-penalizing long ones, and often conflicting with expert judgments. Building on our Multi-Range framework [GladkoffEtAl2024], we propose a calibrated non-linear scoring model that better reflects perceived translation quality across variable-length samples. Empirical evidence from three large-scale production evaluation settings indicates that acceptable error counts grow logarithmically, rather than linearly, with sample size. This trend aligns with psychophysical and cognitive accounts (e.g., the Weber–Fechner law and Cognitive Load Theory), which suggest diminishing marginal perceptual impact of errors alongside increasing cumulative cognitive burden. We introduce a two-parameter tolerance model anchored to a reference point and calibrated from two additional tolerance points via a one-dimensional root-finding procedure. The resulting formulation provides a closed-form range in which the linear approximation remains within $\pm 20\%$ relative error, and integrates into existing MQM workflows with minimal changes by replacing constant tolerance with a length-dependent function. The proposed method improves interpretability, fairness, and inter-rater reliability across both human and AI-generated translations, supporting more scalable and perceptually grounded translation quality evaluation.
@@
Translation quality evaluation, MQM, Human evaluation, Non-linear scoring, Inter-rater reliability
@@ 1 Introduction
Analytic Translation Quality Evaluation (TQE) is widely regarded as a high-precision approach to assessing translation quality. It is based on systematic annotation of linguistic and technical errors using a predefined taxonomy, most commonly derived from the Multidimensional Quality Metrics (MQM) framework [mqm2014, GladkoffEtAl2024], and supported by structured annotator training. Once annotations have been produced, a scoring procedure is applied to convert error patterns into a quantitative quality estimate.
In operational settings, most scoring schemes adopt an implicit linear scaling assumption: the number of permissible errors is expected to increase proportionally with the word count of the evaluated sample. In practice, this typically involves normalizing error counts to a reference sample size (e.g., 1,000 or 2,000 words) and comparing them against a user-defined threshold, such as allowing no more than five minor errors per 1,000 words. The resulting normalized error density is then used either to make an acceptability decision (pass/fail) or to derive a continuous quality score. [^1] Logrus Global Statistics https://logrusglobal.substack.com/archive.
While such linear normalization can be adequate for sample lengths near the calibration reference, it introduces systematic distortions when applied to substantially shorter or longer texts [gladkoff2021]. Practitioners and researchers have repeatedly reported mismatches between linear penalty-based scoring and holistic judgments of translation quality. For instance, Waddington’s comparative studies [waddington1999, waddington2001] show that additive points-off models alone correlate poorly with holistic evaluations, whereas combined analytic–holistic approaches can improve agreement. Similar effects have been documented in related evaluation domains. In speech and video quality assessment, international standards observe saturation effects in opinion scores and recommend comparison-based designs and non-linear mappings (typically logistic) between objective indices and Mean Opinion Scores (MOS) [ituP800, streijl2016, vqeglogistic]. Within translation quality assurance practice, non-linearity is also reflected through severity weighting and the treatment of critical “showstopper” errors, where a single major defect may outweigh many minor ones [mqm2014, saeJ2450]. Large-scale machine translation evaluations further show that expert, context-aware severity judgments yield rankings that diverge from simple averaging, providing additional evidence for non-additive scoring [freitag2021]. Related patterns are well established in usability engineering, where a single catastrophic defect can dominate acceptability decisions [nielsen1994]. Collectively, these findings suggest that perceived translation quality is not well characterized by linear accumulation of penalties.
Despite these known limitations, linear error scaling remains common in industry practice, partly due to institutional convention and partly due to the tendency to apply linear extrapolation in numerical reasoning. Importantly, when evaluators assess translations without explicitly performing normalization, their judgments often deviate from what a linear model would predict. [^2] We provide examples of institutional user reports in Appendix A.
This motivates the research questions addressed in this work:
RQ1: How does the maximum number of tolerable MQM errors change as a function of evaluation sample length in real-world translation quality assessment settings?
RQ2: How can this relationship be operationalized as a calibrated non-linear scoring model that remains interpretable, fair across sample lengths, and compatible with existing MQM-based evaluation workflows?
To answer these questions, we extend our earlier Multi-Range TQE framework [GladkoffEtAl2024] and argue that tolerance for errors increases with sample length, but at a diminishing rate; equivalently, the acceptable error rate per 1,000 words decreases as text length increases. We model this behavior using a logarithmic tolerance function and derive a practical calibration procedure that can be integrated into existing MQM workflows with minimal modification.
Contributions.
This paper makes the following contributions:
  • We provide empirical evidence from three large-scale production evaluation settings addressing RQ1, showing that acceptable error counts grow sublinearly with sample size, consistent with a logarithmic dependence.
  • We link the observed scaling behavior to established psychophysical and cognitive theories, including the Weber–Fechner law and Cognitive Load Theory, thereby motivating a principled non-linear tolerance function.
  • We propose a calibrated two-parameter non-linear tolerance model (RQ2) that can be fit from two tolerance points via a one-dimensional root-finding procedure, yielding a practical scoring function for variable-length evaluation.
  • We show that the proposed model can be integrated into existing MQM-based analytic TQE workflows with minimal modification by replacing constant error tolerance with a length-dependent tolerance function, improving interpretability and supporting more consistent evaluation across sample sizes.
The remainder of this paper is structured as follows. Section 2 reviews analytic TQE and MQM-based scoring. Section 3 presents empirical evidence for non-linear tolerance growth. Section 4 discusses psychophysical and cognitive foundations. Section 5 introduces the proposed non-linear scoring model. Sections 6 and 7 discuss practical implications and provide guidance for constructing a non-linear scorecard. Section 8 concludes with limitations and directions for future work.
@@ 2.1 Analytic TQE and MQM
Analytic Translation Quality Evaluation (TQE) is a structured approach to assessing translation output by identifying, categorizing, and recording observable errors with respect to the source text and relevant specifications. In contrast to holistic evaluation methods, which produce a single overall impression of quality, analytic TQE requires systematic annotation of discrete error instances, each assigned to a predefined error type and (typically) a severity level.
The most widely adopted framework for such structured error annotation is the Multidimensional Quality Metrics (MQM) standard [mqm2014]. MQM defines a hierarchical, extensible typology with high-level error dimensions such as Accuracy and Linguistic Conventions (formerly Fluency), and more fine-grained categories including Mistranslation, Omission, Grammar, and Spelling. In practice, many industry evaluation schemes implement MQM directly or adopt customized subsets tailored to specific content types, client requirements, or quality objectives.
Following annotation, a scoring procedure is applied to an evaluation sample—a defined portion of a translation job—to convert the set of annotated errors into either (i) a numeric quality score or (ii) an acceptability decision (pass/fail). MQM-based evaluation therefore consists of two complementary components: (i) an error typology specifying which error categories and severity levels are annotated, and (ii) a scoring model that maps these annotations to severity-weighted penalty points and an overall score. For comparability across texts of different lengths, scores are typically expressed relative to a fixed reference sample size, most commonly 1,000–2,000 words. [^3] https://themqm.org/.
@@ 2.2 Linear Scoring Models and Their Limits
The dominant scoring approach in analytic TQE is based on linear length normalization. Under this assumption, the severity-weighted penalty total (APT, Absolute Penalty Total; i.e., the sum of penalty points assigned to annotated errors) is scaled proportionally to a reference word count so that results can be compared across evaluation samples of different sizes. For example, a tolerance of five minor errors (one point each) per 1,000 words corresponds to 2.5 penalty points in a 500-word sample and 10 penalty points in a 2,000-word sample. If major errors receive a $5\times$ severity multiplier, then a single major error consumes the full five-point allowance for a 1,000-word sample.
Although proportional scaling provides a simple and operationally convenient normalization rule, it does not fully capture how translation quality is perceived in practice. Empirical and practitioner-oriented work suggests that evaluators apply stricter expectations to very short texts, where individual errors are salient, while showing greater tolerance for longer passages [waddington1999, waddington2001, mqm2014, saeJ2450]. At the same time, tolerance does not increase indefinitely with length: as documents grow, error accumulation and compounding effects contribute to declining acceptability. For instance, a single major error may be sufficient to fail a one-page sample, yet a seven-page sample would not typically be judged acceptable with seven major errors, even though proportional scaling would imply this threshold. Such judgments indicate that the relationship between sample length and acceptable error count is fundamentally non-linear.
@@
Building on the Multi-Range framework proposed in [GladkoffEtAl2024], we distinguish three regimes of evaluation sample size:
  • Micro-range (<250<250 words): For very short samples, analytic scoring exhibits high sampling uncertainty. As shown in [gladkoff2021], estimates derived from such brief segments yield wide confidence intervals, limiting their interpretability for operational decision-making.
    This limitation is also relevant for sentence- or segment-level Quality Estimation (QE). Recent industry observations suggest that fine-grained QE predictions may not consistently reflect document-level acceptability, particularly when critical errors occur sparsely or depend on broader context [baneATA2025, gladkoff2025substack]. These findings motivate treating micro-range scoring as a distinct regime in which statistical uncertainty dominates, and where Statistical Quality Control (SQC) methods can provide a more risk-aware interpretation of limited observations.
@@
  • Meso-range ($\approx 250\!-\!3{,}000$ words): Within this range, scoring based on a reference-length normalization can be operationally effective. In particular, a linearly calibrated scoring model anchored to a user-defined tolerance in severity-weighted penalty points (APT) at a reference word count is commonly used.
@@
  • Macro-range (>3,000>3{,}000 words): For longer samples, human judgments increasingly diverge from linear extrapolation calibrated in the meso-range. In this regime, a non-linear tolerance function—in particular, a logarithmic scaling curve—is better suited to model perceived quality across both meso- and macro-range texts without requiring multiple piecewise linear approximations.
@@
The present study focuses on the meso- and macro-range, where analytic measurements are statistically more stable and practically meaningful. Here, “statistically reliable” refers to estimates with sufficiently narrow confidence intervals, i.e., sampling-uncertainty bounds that shrink as the evaluated sample size increases. [^4] For intuition, a simple 95% Wald interval for a proportion is $p\pm 1.96\sqrt{p(1-p)/n}$ [gladkoff2021]. For small samples or extreme proportions, this estimate is known to undercover; a Wilson (score) interval—or the closely related Agresti–Coull “plus-four” approximation—provides improved coverage [AgrestiCoull1998]. Wilson 95% CI: $\frac{\hat{p}+\frac{z^{2}}{2n}\pm z\sqrt{\frac{\hat{p}(1-\hat{p})}{n}+\frac{z^{2}}{4n^{2}}}}{1+\frac{z^{2}}{n}},$ with z=1.96z\!=\!1.96; the Agresti–Coull “plus-four” uses $\tilde{p}=(x+\tfrac{z^{2}}{2})/(n+z^{2})$ and $\tilde{p}\pm z\sqrt{\tfrac{\tilde{p}(1-\tilde{p})}{n+z^{2}}}$ [AgrestiCoull1998].
@@
Applying linear scoring indiscriminately across all sample sizes often leads to systematic friction in evaluation practice. Users and linguists may need to override, reinterpret, or replace scores, or maintain multiple length-specific scoring rules for short and long samples. A common workaround is to restrict evaluation to a fixed 1,000-word segment; however, this strategy is operationally restrictive and does not address the underlying mismatch between linear scaling and perceived quality. Moreover, because confidence interval width decreases approximately as $1/\sqrt{n}$—with an additional finite-population correction when the evaluated sample constitutes a substantial fraction of the document—variation in sampled word count can materially affect estimation uncertainty. [^5] When the sample is a large fraction of the document, apply the finite-population factor $\sqrt{(N-n)/(N-1)}$ [gladkoff2021] compared to the simple 95% Wald form of $p\pm 1.96\sqrt{p(1-p)/n}$.
The persistence of linear scoring practices, despite these limitations, can be attributed in part to the lack of simple, theoretically grounded alternatives that can be deployed within established MQM workflows. The present work addresses this gap by proposing a calibrated non-linear scoring model, supported by both empirical evidence and theoretical motivation drawn from psychophysics and cognitive-load research.
@@ 3 Empirical Evidence of Non-Linearity
To examine how perceived translation quality varies with evaluation sample length, we conducted an empirical elicitation study with Quality Managers from three large institutional translation programs (anonymized). Each organization operates a high-volume multilingual workflow and applies mature, metric-driven TQE procedures to support routine quality monitoring and acceptance decisions.
@@ 3.1 Survey Design
We reused the extended calibration questionnaire introduced in [GladkoffEtAl2024]. The instrument elicits respondents’ tolerance judgments for multiple sample sizes while explicitly discouraging linear extrapolation, thereby enabling the recovery of latent non-linear tolerance functions.
The questionnaire was designed to capture Quality Managers’ intuitive notion of error tolerance, operationalized as the maximum number of severity-weighted penalty points (APT) that would still be considered acceptable for a given sample size. The key methodological challenge was controlling for respondents’ prior exposure to standard length-normalized scoring schemes, which can lead professionals to default to proportional mental scaling. To mitigate such anchoring effects, respondents were instructed not to compute tolerances from existing formulas or reference thresholds, but instead to report what their organization would consider acceptable based on professional judgment and operational experience.
The questionnaire contained the following prompts:
Intuitive Error Tolerance Questions Intuitively, how many minor errors would you accept in a 1-page sample? Without doing any math or using a scoring model, what number of minor errors feels acceptable for 2 pages? For a 3-page text, how many minor errors seem tolerable—based on your experience? Do not calculate—just estimate: how many minor errors would you allow in 4 pages? Based on your intuition, how many minor errors would still be acceptable in a 5-page sample? Without applying your formal model, what number of minor errors seems acceptable in 10 pages? Without referencing formulas, what minor error count would feel acceptable for 20 pages?
@@
To reduce anchoring and promote independent judgments across lengths, page sizes were presented in randomized order and prior responses were not displayed. Two page sizes were repeated at the end of the questionnaire to probe intra-rater consistency. Respondents were again explicitly instructed not to prorate from known thresholds and to rely on professional judgment rather than calculation.
The elicitation focused on minor errors for two reasons. First, higher-severity errors (major and critical) occur relatively infrequently, which makes them less suitable for robust quantitative elicitation at the sample sizes considered. Second, their perceptual and operational impact differs qualitatively, as a small number of high-severity defects may dominate acceptability decisions. In industry practice, severity-weighted scoring allows the effect of any error to be expressed as an equivalent number of minor errors, making minor-error tolerance a practical proxy for overall tolerance modeling.
Respondents expressed sample length either in pages or in word counts. For comparability across institutions, all answers were normalized to pages using a conversion factor of 250 words per page.